# Topology and Diffeology via Metric-like Functions

### v2

@@ 5. Diffeology defined by $d_{r}$
By the inequality $d_{r^{\prime}}(p(t_{0}),p(t))\geq\varepsilon$, it follows that
$(p(t_{0})+v)\cap{r^{\prime}}D\neq p(t)\cap{r^{\prime}}D$
for the same $v$. Therefore, we obtain
$d_{r^{\prime}}(p(t_{0}),p(t))\geq\min_{\tau\in\mathscr{P}}(\mathop{\mathrm{diam}}{\tau})-\varepsilon>\varepsilon.$
On the other hand, since $t\in\overline{A}_{r^{\prime}}$, there exists a sequence $\{s_{n}\}_{n\in\mathbb{Z}_{\geq 0}}$ converging to $t$ such that
$d_{r^{\prime}}(p(t_{0}),p(s_{n}))<\varepsilon.$
This contradicts the continuity of the map $d_{r^{\prime}}(p(t_{0}),p(\cdot))\colon X\to\mathbb{R}$ at the point $t$. Thus, we have $\overline{A_{r^{\prime}}}\setminus A_{r^{\prime}}\neq\emptyset$, which implies that $A_{r^{\prime}}$ is closed in $B_{\delta}{(t_{0})}$. Therefore, we obtain $A_{r^{\prime}}=B_{\delta}{(t_{0})}$.
By the above argument, for each $t\in B_{\delta}{(t_{0})}$, there exists a unique $u_{t}\in\mathbb{R}^{d}$ such that $|u_{t}|<\varepsilon$ and
@@
for each $r^{\prime}>0$. This $u_{t}$ satisfies
$p(t_{0})+u_{t}=p(t)$
for each $t\in B_{\delta}{(t_{0})}$. Therefore, we have a unique map $u\colon B_{\delta}{(t_{0})}\to\mathbb{R}^{d};t\mapsto u_{t}$ such that $p(t_{0})+u(t)=p(t)$ for each $t\in B_{\delta}{(t_{0})}$. We now show that this map $u$ is smooth.
First, we prove the continuity of $u$ at the point $t_{0}\in B_{\delta}{(t_{0})}$. It has already been proved that the map $d_{r}(p(t_{0}),p(\cdot))\colon B_{\delta}{(t_{0})}\to\mathbb{R}$ is continuous. Since $d_{r}(p(t_{0}),p(t))<\varepsilon$ for each $t\in B_{\delta}{(t_{0})}$, it follows that
$d_{r}(p(t_{0}),p(t))=|u(t_{0})-u(t)|.$
Therefore, for any $\varepsilon^{\prime}>0$, there exists $\delta^{\prime}>0$ such that for each $t\in B_{\delta^{\prime}}{(t_{0})}$,
$|u(t_{0})-u(t)|=|u(t_{0})-u(t)|-|u(t_{0})-u(t_{0})|<\varepsilon^{\prime},$
which implies that $u$ is continuous at $t_{0}$.
Next, we prove the smoothness of $u$ at the point $t_{0}\in B_{\delta}{(t_{0})}$. If $p(t)=p(t_{0})$ holds for all $t\in B_{\delta}{(t_{0})}$, then $u$ is constant on $B_{\delta}{(t_{0})}$, and in particular, smooth on $B_{\delta}{(t_{0})}$. Otherwise, there exists $t_{1}\in B_{\delta}{(t_{0})}\setminus p^{-1}(p(t_{0}))$. By repeating this process, we can find a finite sequence $t_{1},\dots,t_{n}\in B_{\delta}{(t_{0})}$ such that
$\{u(t)\mid t\in B_{\delta}{(t_{0})}\}\subset\mathop{\mathrm{Span}}\{\overrightarrow{u(t_{0})u(t_{i})}\mid i=1,\dots,n\}$
and the set of vectors $\left\{\overrightarrow{u(t_{0})u(t_{i})}\right\}_{i=1}^{n}$ is linearly independent. Thus, we can restrict the codomain of $u$ to $\mathop{\mathrm{Span}}\left\{\overrightarrow{u(t_{0})u(t_{i})}\mid i=1,\dots,n\right\}\cong\mathbb{R}^{n}$. We define the map
$f\colon B_{\delta}{(t_{0})}\times\mathbb{R}^{n}\to\mathbb{R}^{n};(t,x)\mapsto(|x-u(t_{i})|^{2}-d_{r}(p(t_{i}),p(t))^{2})_{i=1}^{n}.$
Since the map $d_{r}(p(t_{i}),p(\cdot))$ is smooth at $t_{0}$ for each $t_{i}$, the map $f$ is smooth around $(t_{0},u(t_{0}))$. By the linear independence of $\left\{\overrightarrow{u(t_{0})u(t_{i})}\right\}_{i=1}^{n}$, we have
$\left.\det\left({\frac{\partial f}{\partial x}}(t_{0},u(t_{0}))\right)=\det{\begin{pmatrix}2x-2u(t_{1})\\ \vdots\\ 2x-2u(t_{n})\end{pmatrix}}\right|_{t=t_{0},x=u(t_{0})}\neq 0.$
Thus, by the implicit function theorem, there exist open sets $U\subset B_{\delta}{(t_{0})}$ and $V\subset\mathbb{R}^{n}$ with $t_{0}\in U$ and $u(t_{0})\in V$, and a smooth map $g\colon U\to V$ such that for all $(t,x)\in U\times V$,
@@
6. Relating topologies and diffeologies
@@
In [11], a diffeology on $\Psi_{d}(\mathbb{R}^{N})$ is defined as follows: a parametrization $p\colon U_{p}\to\Psi_{d}(\mathbb{R}^{N})$ is a plot if
@@
is a closed submanifold of $\mathbb{R}^{N}$, and the projection
$\hat{p}\colon\Gamma_{p}\to U_{p},\quad(t,x)\mapsto t$
@@
It is natural to ask how this diffeology relates to the one induced by by $\{d_{r}\}_{r>0}$ (see Proposition 5.6). At least, the following fact holds in general.
Lemma 6.1.
Let $(X,\{d_{r}\}_{r>0})$ satisfy the weaker $\{d_{r}\}$-axioms in Definition 4.4. If a map $p\colon U_{p}\to X$ is a plot of $X$ as defined in Proposition 5.6, then $p$ is also continuous.
@@
Proposition 6.2.
Let $(X,\{d_{r}\}_{r>0})$ satisfy the weaker $\{d_{r}\}$-axioms in Definition 4.4. The $D$-topology of the diffeology on $X$ induced by the family $\{d_{r}\}_{r>0}$ (see Proposition 5.6) is finer than the topology on $X$ induced by the same family (see Proposition 4.6).
@@
Let $x\in X$, $r>0$, and $\varepsilon>0$ be arbitrary. For each plot $p\colon U_{p}\to X$ of $X$, the preimage $p^{-1}(\mathcal{U}_{r,\varepsilon}(x))$ is open in $U_{p}$ because $p$ is continuous (Lemma 6.1). Therefore, $\mathcal{U}_{r,\varepsilon}(x)$ is $D$-open. ∎

### v3

@@ 5. Diffeology defined by $d_{r}$
Since $t\notin A_{r^{\prime}}$, the same $v$ does not satisfy
$(p(t_{0})+v)\cap{r^{\prime}}D=p(t)\cap{r^{\prime}}D.$
If $d_{r^{\prime}}(p(t_{0}),p(t))<\lambda_{T}$, then there exists $u\in\mathbb{R}^{d}$ with $|u|<\lambda_{T}$ such that $(p(t_{0})+u)\cap{r^{\prime}}D=p(t)\cap{r^{\prime}}D$. Restricting to rD{r}D, we also have $(p(t_{0})+u)\cap{r}D=p(t)\cap{r}D$, and since $|v|<\varepsilon<\lambda_{T}$, the defining property of $\lambda_{T}$ implies $u=v$, a contradiction. Hence,
$d_{r^{\prime}}(p(t_{0}),p(t))\geq\lambda_{T}>\varepsilon.$
On the other hand, since $t\in\overline{A}_{r^{\prime}}$, there exists a sequence $\{s_{n}\}_{n\in\mathbb{Z}_{\geq 0}}$ converging to $t$ such that $d_{r^{\prime}}(p(t_{0}),p(s_{n}))<\varepsilon$ for all $n$. This contradicts the continuity of the map $t\mapstochar\rightarrow d_{r^{\prime}}(p(t_{0}),p(t))$ at the point $t$. Thus, $A_{r^{\prime}}$ is closed in $B_{\delta^{\prime}}{(t_{0})}$. Therefore, we obtain $A_{r^{\prime}}=B_{\delta^{\prime}}{(t_{0})}$.
By the above argument, for each $t\in B_{\delta^{\prime}}{(t_{0})}$ and each $r^{\prime}>0$, there exists $u\in\mathbb{R}^{d}$ such that $|u|<\varepsilon$ and
$(p(t_{0})+u)\cap{r^{\prime}}D=p(t)\cap{r^{\prime}}D.$
By the defining property of $\lambda_{T}$, such a vector $u$ is unique, and hence independent of $r^{\prime}$. Therefore, for each $t\in B_{\delta^{\prime}}{(t_{0})}$, there exists a unique $u_{t}\in\mathbb{R}^{d}$ such that $|u_{t}|<\varepsilon$ and
@@
for every $r^{\prime}>0$. In particular, $p(t_{0})+u_{t}=p(t)$ for each $t\in B_{\delta^{\prime}}{(t_{0})}$. Thus, we obtain a unique map $u\colon B_{\delta^{\prime}}{(t_{0})}\rightarrow\mathbb{R}^{d};\ t\mapstochar\rightarrow u_{t}$ such that $p(t_{0})+u(t)=p(t)$ for each $t\in B_{\delta^{\prime}}{(t_{0})}$. We now show that this map $u$ is smooth.
First, we prove the continuity of $u$ at the point $t_{0}\in B_{\delta^{\prime}}{(t_{0})}$. It has already been proved that the map $d_{r}(p(t_{0}),p(\cdot))\colon B_{\delta^{\prime}}{(t_{0})}\rightarrow\mathbb{R}$ is continuous. Since $d_{r}(p(t_{0}),p(t))=|u(t_{0})-u(t)|$ for each $t\in B_{\delta^{\prime}}{(t_{0})}$, for any $\varepsilon^{\prime}>0$, there exists $\eta>0$ such that for each $t\in B_{\eta}{(t_{0})}$,
$|u(t)-u(t_{0})|=\bigl|d_{r}(p(t_{0}),p(t))-d_{r}(p(t_{0}),p(t_{0}))\bigr|<\varepsilon^{\prime}.$
Hence, $u$ is continuous at $t_{0}$.
Next, we prove the smoothness of $u$ at the point $t_{0}\in B_{\delta^{\prime}}{(t_{0})}$. If $p(t)=p(t_{0})$ holds for all $t\in B_{\delta^{\prime}}{(t_{0})}$, then $u$ is constant on $B_{\delta^{\prime}}{(t_{0})}$, and in particular, smooth on $B_{\delta^{\prime}}{(t_{0})}$. Otherwise, there exists $t_{1}\in B_{\delta^{\prime}}{(t_{0})}\setminus p^{-1}(p(t_{0}))$. By repeating this process, we can find a finite sequence $t_{1},\dots,t_{n}\in B_{\delta^{\prime}}{(t_{0})}$ such that
$\{u(t)\mid t\in B_{\delta^{\prime}}{(t_{0})}\}\subset\mathop{\mathrm{Span}}\{\overrightarrow{u(t_{0})u(t_{i})}\mid i=1,\dots,n\}$
and the set of vectors $\left\{\overrightarrow{u(t_{0})u(t_{i})}\right\}_{i=1}^{n}$ is linearly independent. Thus, we can restrict the codomain of $u$ to $\mathop{\mathrm{Span}}\left\{\overrightarrow{u(t_{0})u(t_{i})}\mid i=1,\dots,n\right\}\cong\mathbb{R}^{n}$. Note that $|u_{t}|<\varepsilon<\lambda_{T}/2$ for each $t\in B_{\delta^{\prime}}{(t_{0})}$, so for any $t,t_{i}\in B_{\delta^{\prime}}{(t_{0})}$, we have $|u(t)-u(t_{i})|\leq|u(t)|+|u(t_{i})|<2\varepsilon<\lambda_{T}$. Therefore, the same uniqueness argument gives $d_{r}(p(t_{i}),p(t))=|u(t_{i})-u(t)|$ for all $t,t_{i}\in B_{\delta^{\prime}}{(t_{0})}$. We define the map
$f\colon B_{\delta^{\prime}}{(t_{0})}\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{n};(t,x)\mapstochar\rightarrow(|x-u(t_{i})|^{2}-d_{r}(p(t_{i}),p(t))^{2})_{i=1}^{n}.$
In particular, $f(t_{0},u(t_{0}))=0$ and $f(t,u(t))=0$ for all $t\in B_{\delta^{\prime}}{(t_{0})}$. Since the map $d_{r}(p(t_{i}),p(\cdot))$ is smooth at $t_{0}$ for each $t_{i}$, the map $f$ is smooth around $(t_{0},u(t_{0}))$. By the linear independence of $\left\{\overrightarrow{u(t_{0})u(t_{i})}\right\}_{i=1}^{n}$, we have
$\left.\det\left({\frac{\partial f}{\partial x}}(t_{0},u(t_{0}))\right)=\det{\begin{pmatrix}{}^{t}(2x-2u(t_{1}))\\ \vdots\\ {}^{t}(2x-2u(t_{n}))\end{pmatrix}}\right|_{t=t_{0},x=u(t_{0})}\neq 0.$
Thus, by the implicit function theorem, there exist open sets $U\subset B_{\delta^{\prime}}{(t_{0})}$ and $V\subset\mathbb{R}^{n}$ with $t_{0}\in U$ and $u(t_{0})\in V$, and a smooth map $g\colon U\rightarrow V$ such that for all $(t,x)\in U\times V$,
@@
Appendix A Relating topologies and diffeologies
@@
In [MR4821355], a diffeology on $\Psi_{d}(\mathbb{R}^{N})$ is defined as follows: a parametrization $p\colon U_{p}\rightarrow\Psi_{d}(\mathbb{R}^{N})$ is a plot if
@@
is a closed submanifold of $U_{p}\times\mathbb{R}^{N}$, and the projection
$\hat{p}\colon\Gamma_{p}\rightarrow U_{p},\quad(t,x)\mapstochar\rightarrow t$
@@
It is natural to ask how this diffeology relates to the one induced by $\{d_{r}\}_{r>0}$ (see Proposition 5.6). At least, the following fact holds in general.
Lemma A.1.
Let $(X,\{d_{r}\}_{r>0})$ satisfy the weaker $\{d_{r}\}$-axioms in Definition 4.4. If a map $p\colon U_{p}\rightarrow X$ is a plot of $X$ as defined in Proposition 5.6, then $p$ is also continuous.
@@
Proposition A.2.
Let $(X,\{d_{r}\}_{r>0})$ satisfy the weaker $\{d_{r}\}$-axioms in Definition 4.4. The $D$-topology of the diffeology on $X$ induced by the family $\{d_{r}\}_{r>0}$ (see Proposition 5.6) is finer than the topology on $X$ induced by the same family (see Definition 2.3).
@@
Let $x\in X$, $r>0$, and $\varepsilon>0$ be arbitrary. For each plot $p\colon U_{p}\rightarrow X$ of $X$, the preimage $p^{-1}(\mathcal{U}_{r,\varepsilon}(x))$ is open in $U_{p}$ because $p$ is continuous (Lemma A.1). Therefore, $\mathcal{U}_{r,\varepsilon}(x)$ is $D$-open. ∎